OpenStudy (anonymous):

What is the slope-intercept form of this equation? 8x + 6y = 12 A. y= -4/3x + 2 B. y= -4/3x + 2 C. y= 4/3x + 6 D. y= 2/3x + 2

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@KiaKat

OpenStudy (igreen):

First subtract 8x to both sides

OpenStudy (anonymous):

- 2y + -4 = 8x?

OpenStudy (igreen):

Um, no.. 8x + 6y = 12 Subtract 8x to both sides, leave everything else the same.

OpenStudy (anonymous):

0x + 2y? im so confused

OpenStudy (igreen):

No, subtracting will just move it to the other side: 8x + 6y = 12 6y = -8x + 12 Now divide 6 to both sides, what's 6 / 6?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

?

OpenStudy (igreen):

Yes, that leaves us with just 'y' Can you simplify -8/6?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

?

OpenStudy (anonymous):

would it happen to be c?

OpenStudy (anonymous):

it would be -4/3

OpenStudy (anonymous):

@KiaKat

OpenStudy (igreen):

Yes, now divide 12 / 2

OpenStudy (igreen):

@StonerWill

OpenStudy (anonymous):

6

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Oops sorry, I meant 12 / 6

OpenStudy (anonymous):

2

OpenStudy (anonymous):

there for A.?

OpenStudy (igreen):

Yep, so we have: \(y = -\dfrac{4}{3}x + 2\)

OpenStudy (igreen):

I don't know, A and B look the same.

OpenStudy (anonymous):

thats my error but thankss(:

OpenStudy (igreen):

No problem :)